OpenStudy (anonymous):

A number is decreased by seven. Then, the new number is multiplied by 3 to get an answer of -9. What is the original number? -30 30 4 -4

OpenStudy (anonymous):

its 4

OpenStudy (potatoes.ramu):

Let the number be x (x-7)*3 = -9 This implies that, 3x - 21 = -9 Thus, 3x = 12 So, x = 4.

OpenStudy (anonymous):

let the original number be "x" (x-7)*3 = -9 so, expand the bracket using Bodmas (bidmas) = Brackets , Other (indices), division, Multiplication, addition and subtraction. Do this in order starting with brackets, then, other, then multiplication and so on... 3x - 21 = -9 so, 3x = 12 x = 12/3 therefore, "x = 4"